# “Cold, Calculated, and Condescending”: How AI Identifies and Explains Ableism Compared to Disabled People

### v2

@@ 6. Limitations and Future Work
This study focuses on ableist speech in text-based online content and does not extend to other forms of AI, such as generative images (Glazko et al., 2023; Tevissen, 2024). A primary limitation of this work is our dataset size—200 comments, sourced from reported online ableism by PwD, cannot fully represent the spectrum of ableist language. Similarly, with 130 PwD participants, our findings are not universally representative of all disabled individuals. Larger scale studies are required to explore relationships between participant demographics and ableist harm. Despite these constraints, this research offers a critical first step in examining how AI models interact with ableist speech. This work goes beyond model upgrades and variations in ratings, putting disabled individuals at the forefront to critically evaluate AI output, and laying the groundwork for understanding how AI models detect and explain disability bias.
@@ 7. Ethical Considerations
This study is situated in a context within which, without proper safeguards, research activities could cause unintended harm to research participants. For example, people with disabilities might experience distress when labeling posts that contain ableist hate. Following best practices from Chen et al. (Chen et al., 2022) and Bellini et al. (Bellini et al., 2024) on working with vulnerable communities, we took a number of steps to conduct our research ethically and responsibly. For example, we designed our surveys conscientiously, and deflected personal impact of harm by asking questions in third person (e.g., ”A person with a disability goes on social media and reads the following comment.”). We further ensured a lower cognitive load by restricting the length and type of questions displayed (i.e., no grid questions) (Wilson et al., 2013; Brosnan et al., 2021). We also allowed participants to skip questions or exit the study at any point. While the surveys were expected to take 15 minutes, participants were given up to an hour to complete them. They were fairly compensated up to \$10 via Prolific, equating to an estimated \$40/hour for completing a 15-minute survey. Additionally, we provided mental health resources to help participants manage emotional distress and shared contact details for any study-related concerns.
All aspects of this research were reviewed and approved by the IRB at our institution. Participants were informed of potential risks and gave their informed consent to share anonymized research data. During data collection and release, no personally identifiable information (PII) was collected from participants or distributed to LLMs.
@@ A.2. LLM Prompt
This prompt was used to correspond with all LLMs: GPT-4, Gemini, Claude, and Llama.
@@
Ableist Comments Non-Ableist Comments
Pairwise Groups Z (P-Value) Effect Size Z (P-Value) Effect Size
Perspective – PwD –14.20 (0.000) *** 1.004 –1.52 (1.000)
AzureAI – PwD – 11.26 (0.000) *** 0.796 – 3.48 (0.017) * 0.246
OpenAI – PwD – 9.52 (0.000) *** 0.673 – 1.16 (1.000)
GPT-4 – PwD – 3.88 (0.000) *** 0.274 1.79 (1.000)
Gemini – PwD – 8.03 (0.000) *** 0.568 – 1.44 (1.000)
Claude – PwD – 2.83 (0.000) *** 0.210 – 0.09 (1.000)
Llama – PwD – 0.60 (0.545) 1.03 (1.000)
AzureAI – Non-PwD – 6.75 (0.000) *** 0.478 – 2.78 (0.153)
OpenAI – Non-PwD – 4.81 (0.000) *** 0.340 0.18 (1.000)
Perspective – Non-PwD – 7.59 (0.000) *** 0.537 – 0.94 (1.000)
GPT-4 – Non-PwD 0.63 (1.000) 1.92 (1.000)
Gemini – Non-PwD – 2.73 (0.307) * – 0.88 (1.000)
Claude – Non-PwD 1.47 (0.561) – 0.29 (1.000)
Llama – Non-PwD 3.28 (0.008) ** 0.232 – 1.26 (1.000)
PwD – Non-PwD 3.42 (0.000) *** 0.242 0.37 (1.000)
Table 2. A.3.1 Wilcoxon-Signed Rank Test Results for Toxicity Ratings, between AI Models, PwD, and Non-PwD
Figure 4. A.3.2 Spearman Correlation Scores for Toxicity Ratings between AI Models, PwD, and Non-PwD
Figure 5. A.3.3 Box Plots for Implicit and Explicit Toxicity Ratings
Ableist Comments Non-Ableist Comments
Pairwise Groups Z (P-Value) Effect Size Z (P-Value) Effect Size
Claude – Gemini 11.57 (0.000) *** 0.818 4.23 (0.000) *** 0.299
Claude – GPT-4 2.44 (0.005) ** 0.172 1.77 (0.680)
Gemini – GPT-4 – 9.13 (0.000) *** 0.646 – 2.45 (0.170) 0.330
Gemini – Llama – 8.23 (0.000) *** 0.582 – 3.62 (0.004) ** 0.256
Claude – Llama 3.34 (0.000) *** 0.236 0.61 (1.000)
GPT-4 – Llama 0.897 (0.370) - 1.17 (1.000)
GPT-4 – PwD 2.23 (0.000) *** 0.158 1.16 (0.980)
Gemini – PwD – 6.90 (0.000) *** 0.488 – 1.28 (1.000)
Claude – PwD 4.67 (0.000) *** 0.330 2.939 (0.042) ** 0.208
Llama – PwD 1.33 (0.37) 2.361 (0.02) * 0.167
GPT-4 – Non-PwD 4.66 (0.000) *** 0.329 0.718 (1.000)
Gemini – Non-PwD – 2.72 (0.032) *** 0.19 – 1.49 (0.957)
Claude – Non-PwD 6.63 (0.000) *** 0.469 2.32 (0.204)
Llama – Non-PwD 3.93 (0.000) *** 0.278 1.77 (0.615)
PwD – Non-PwD 3.93 (0.000) *** 0.278 0.327 (0.743) **
Table 3. A.4.1 Pairwise Wilcoxon-Signed Rank for Ableism Ratings, between AI Models, PwD, and Non-PwD
Figure 6. A.4.2 Spearman Correlation Scores for Ableism Ratings between LLMs, PwD, and Non-PwD
Figure 7. A.4.3 Box Plots for Implicit and Explicit Ableism Ratings
Figure 8. Density Distributions of Toxicity and Ableism Ratings by LLMs, PwD, and Non-PwD

### v3

@@ 6. Limitations and Future Work
This study focuses on ableist speech in text-based online content and does not extend to other forms of AI, such as generative images (Glazko et al., 2023; Tevissen, 2024). A primary limitation of this work is our dataset size—200 comments, sourced from reported online ableism by PwD, cannot fully represent the spectrum of ableist language. Similarly, with 130 PwD participants, our findings are not universally representative of all disabled individuals. Larger scale studies are required to explore relationships between participant demographics and ableist harm. Despite these constraints, this research offers a critical first step in examining how AI models interact with ableist speech.
This work goes beyond model upgrades and variations in ratings, putting disabled individuals at the forefront to critically evaluate AI output, and laying the groundwork for understanding how AI models detect and explain disability bias.
@@ 7. Ethical Considerations
This study is situated in a context within which, without proper safeguards, research activities could cause unintended harm to research participants. For example, people with disabilities might experience distress when labeling posts that contain ableist hate. Following best practices from Chen et al. (Chen et al., 2022) and Bellini et al. (Bellini et al., 2024) on working with vulnerable communities, we took a number of steps to conduct our research ethically and responsibly. For example, we designed our surveys conscientiously, and deflected personal impact of harm by asking questions in third person (e.g., ”A person with a disability goes on social media and reads the following comment.”). We further ensured a lower cognitive load by restricting the length and type of questions displayed (i.e., no grid questions) (Wilson et al., 2013; Brosnan et al., 2021). We also allowed participants to skip questions or exit the study at any point. While the surveys were expected to take 15 minutes, participants were given up to an hour to complete them. They were fairly compensated up to \$10 via Prolific, equating to an estimated \$40/hour for completing a 15-minute survey. Additionally, we provided mental health resources to help participants manage emotional distress and shared contacts for any study-related concerns.
All aspects of this research were reviewed and approved by the IRB at our institution. Participants were informed of potential risks and gave their informed consent to share anonymized research data. During data collection and release, no personally identifiable information was collected from participants or distributed to LLMs.
Acknowledgements.
We are grateful to the Dennis Washington Foundation, Cornell Center for Social Sciences, Einaudi Center for International Studies, and Google for supporting this work. We thank our participants for their time, their insights, and their vulnerability.
@@ A.2. LLM Prompt
This prompt was used to correspond with all LLMs:
@@
A.3. Statistical Test Results for Toxicity Identification
Wilcoxon-Signed Rank Test Results for Toxicity Ratings, between AI Models, PwD, and Non-PwD.
Pairwise Groups Z (P-Value) Effect Size
Perspective – PwD –14.20 (0.000) *** 1.004
AzureAI – PwD –11.26 (0.000) *** 0.796
OpenAI – PwD –9.52 (0.000) *** 0.673
GPT-4 – PwD –3.88 (0.000) *** 0.274
Gemini – PwD –8.03 (0.000) *** 0.568
Claude – PwD –2.83 (0.000) *** 0.210
Llama – PwD –0.60 (0.545)
AzureAI – Non-PwD –6.75 (0.000) *** 0.478
OpenAI – Non-PwD –4.81 (0.000) *** 0.340
Perspective – Non-PwD –7.59 (0.000) *** 0.537
GPT-4 – Non-PwD 0.63 (1.000)
Gemini – Non-PwD –2.73 (0.307) *
Claude – Non-PwD 1.47 (0.561)
Llama – Non-PwD 3.28 (0.008) ** 0.232
PwD – Non-PwD 3.42 (0.000) *** 0.242
Table 2. A.3.1 Wilcoxon-Signed Rank Test Results for Ableist Comments
Pairwise Groups Z (P-Value) Effect Size
Perspective – PwD –1.52 (1.000)
AzureAI – PwD –3.48 (0.017) * 0.246
OpenAI – PwD –1.16 (1.000)
GPT-4 – PwD 1.79 (1.000)
Gemini – PwD –1.44 (1.000)
Claude – PwD –0.09 (1.000)
Llama – PwD 1.03 (1.000)
AzureAI – Non-PwD –2.78 (0.153)
OpenAI – Non-PwD 0.18 (1.000)
Perspective – Non-PwD –0.94 (1.000)
GPT-4 – Non-PwD 1.92 (1.000)
Gemini – Non-PwD –0.88 (1.000)
Claude – Non-PwD –0.29 (1.000)
Llama – Non-PwD –1.26 (1.000)
PwD – Non-PwD 0.37 (1.000)
Table 3. A.3.2 Wilcoxon-Signed Rank Test Results for Non-Ableist Comments
A.4. Statistical Test Results for Ableism Identification
Pairwise Wilcoxon-Signed Rank for Ableism Ratings, between AI Models, PwD, and Non-PwD.
Pairwise Groups Z (P-Value) Effect Size
Claude – Gemini 11.57 (0.000) *** 0.818
Claude – GPT-4 2.44 (0.005) ** 0.172
Gemini – GPT-4 –9.13 (0.000) *** 0.646
Gemini – Llama –8.23 (0.000) *** 0.582
Claude – Llama 3.34 (0.000) *** 0.236
GPT-4 – Llama 0.897 (0.370)
GPT-4 – PwD 2.23 (0.000) *** 0.158
Gemini – PwD –6.90 (0.000) *** 0.488
Claude – PwD 4.67 (0.000) *** 0.330
Llama – PwD 1.33 (0.37)
GPT-4 – Non-PwD 4.66 (0.000) *** 0.329
Gemini – Non-PwD –2.72 (0.032) *** 0.19
Claude – Non-PwD 6.63 (0.000) *** 0.469
Llama – Non-PwD 3.93 (0.000) *** 0.278
PwD – Non-PwD 3.93 (0.000) *** 0.278
Table 4. A.4.1 Pairwise Wilcoxon-Signed Rank for Ableism Ratings – Ableist Comments
Pairwise Groups Z (P-Value) Effect Size
Claude – Gemini 4.23 (0.000) *** 0.299
Claude – GPT-4 1.77 (0.680)
Gemini – GPT-4 –2.45 (0.170) 0.330
Gemini – Llama –3.62 (0.004) ** 0.256
Claude – Llama 0.61 (1.000)
GPT-4 – Llama -1.17 (1.000)
GPT-4 – PwD 1.16 (0.980)
Gemini – PwD –1.28 (1.000)
Claude – PwD 2.939 (0.042) ** 0.208
Llama – PwD 2.361 (0.02) * 0.167
GPT-4 – Non-PwD 0.718 (1.000)
Gemini – Non-PwD –1.49 (0.957)
Claude – Non-PwD 2.32 (0.204)
Llama – Non-PwD 1.77 (0.615)
PwD – Non-PwD 0.327 (0.743) **
Table 5. A.4.2 Pairwise Wilcoxon-Signed Rank for Ableism Ratings – Non-Ableist Comments